### v1

# Towards Robust Surrogate Models: Benchmarking Machine Learning Approaches to Expediting Phase Field Simulations of Brittle Fracture

Abstract
Data driven approaches have the potential to make modeling complex, nonlinear physical phenomena significantly more computationally tractable. For example, computational modeling of fracture is a core challenge where machine learning techniques have the potential to provide a much needed speedup that would enable progress in areas such as mutli-scale modeling and uncertainty quantification. Currently, phase field modeling (PFM) of fracture is one such approach that offers a convenient variational formulation to model crack nucleation, branching and propagation. To date, machine learning techniques have shown promise in approximating PFM simulations. However, most studies rely on overly simple benchmarks that do not reflect the true complexity of the fracture processes where PFM excels as a method. To address this gap, we introduce a challenging dataset based on PFM simulations designed to benchmark and advance ML methods for fracture modeling. This dataset includes three energy decomposition methods, two boundary conditions, and 1,000 random initial crack configurations for a total of 6,000 simulations. Each sample contains 100 time steps capturing the temporal evolution of the crack field. Alongside this dataset, we also implement and evaluate Physics Informed Neural Networks (PINN), Fourier Neural Operators (FNO) and UNet models as baselines, and explore the impact of ensembling strategies on prediction accuracy. With this combination of our dataset and baseline models drawn from the literature we aim to provide a standardized and challenging benchmark for evaluating machine learning approaches to solid mechanics. Our results highlight both the promise and limitations of popular current models, and demonstrate the utility of this dataset as a testbed for advancing machine learning in fracture mechanics research.
11footnotetext: corresponding author
1 Introduction
In recent years, data-driven modeling has gained significant traction across science and engineering due to its ability to approximate complex physical systems at a fraction of the computational cost of traditional high fidelity simulations [1, 2] for autonomous and non-autonomous systems where the input space varies with time [3]. By extracting patterns from data and learning compact representations, neural network based approximations can offer fast and flexible alternatives to conventional solvers. A wide range of architectures have been proposed to introduce inductive biases relevant to the underlying structure of physical problems. For instance, convolutional neural networks (CNNs) have been applied to structured grids [4] while graph neural networks (GNNs) have been applied to problems defined on unstructured meshes and irregular underlying structures [5, 6]. To better integrate governing physics into the learning process, physics informed neural networks (PINNs) [7] were introduced. These models use neural networks to parameterize the solution of partial differential equations (PDEs) and incorporate the residuals of the governing equations directly into the loss function. PINNs have been applied to a broad range of problems [8, 9, 10, 11, 12], but face scalability challenges as they typically need to be retrained for every change in initial or boundary condition. To overcome these limitations, operator learning approaches have emerged. These methods aim to learn mappings between infinite dimensional input and output function spaces, enabling generalization across different input space parameters such as initial and boundary conditions. Both purely data driven and physics constrained variants have been proposed within this framework achieving state of the art performance across a wide range of scientific and engineering domains [13, 14, 15, 16, 17, 18].
However, despite these advances, neural network based solutions continue to face major challenges. For example, the models mentioned previously, apply the governing physical laws as soft penalties within the loss function rather than enforcing them strictly. As a result, the learned solutions may violate the underlying physics especially in regions with sparse data or high complexity [19, 20, 21, 22]. Additionally, many models suffer from what is known as spectral bias which is the tendency of the model to learn low frequency (smooth) components of the solution first while struggling to capture high frequency features [23, 24]. This makes it challenging for these models to accurately capture regions with sharp gradients, such as discontinuities or localized damage. This behavior is particularly problematic for modeling fracture where small scale features such as cracks play a critical role in determining the system’s overall response. Other persistent issues include difficulty of optimizing the neural networks in a non convex loss landscape [25, 26, 27] and sensitivity to initialization or training dynamics [28, 29, 30]. Together, these limitations make it difficult to develop robust, generalizable and physically faithful models highlighting the need for structured, high fidelity benchmark datasets to systematically study and stress test model capabilities such as the Mechanical MNIST collection [31, 32, 33] which presented accessible benchmark datasets for mechanical engineering inspired from the computer vision research community or other benchmarks that target broader scientific meta modeling benchmarks such as the Well [34] or PDEBench [35].
Inspired by these challenges, we have chosen brittle fracture as the starting point for our dataset. There are many different ways to model brittle fracture [36, 37, 38]. Among these methods, Phase field modeling (PFM) of brittle fracture is a popular modern option that has been explored recently within the ML literature [39, 40]. Notably fracture problems are inherently characterized by discontinuities, sharp gradients, and topological changes such as crack nucleation, branching, and coalescence, which, as mentioned, are particularly difficult for data driven models to capture accurately. At the same time, the coupled governing equations of the PFM provides a well defined and physically grounded framework with intrinsic non convexity [41] making it a natural setting to embed physics into neural networks and rigorously evaluate their robustness and convergence behavior under complex physical settings. PFMs have also been explored in the context of stochastic modeling [41] and uncertainty quantification [42], and are notorious for being computationally expensive [43, 44, 45]. Different attempts have been made by researchers to offer a data driven approximation to the solution of phase field models of brittle fracture [46, 47, 40, 48, 17]. However many of these studies have ultimately evaluated ML models on relatively simple benchmark problems. Without a rigorous exploration of more complex and diverse fracture scenarios, it remains unclear whether these ML models are truly able to capture the subtleties of fracture prediction.
In this work our primary goal is to develop a challenging dataset that enables a deeper exploration of the subtleties of surrogate modeling with a particular focus on phase field modeling of brittle fracture. As can be seen in Fig. 1 this dataset is designed not only to reflect the complex nature of fracture and crack propagation under various boundary conditions and energy decompositions, but also to serve as a robust benchmark for evaluating the performance of data driven surrogate models. Specifically, we implement and assess the performance of several baseline neural network based models including Physics informed neural networks (PINNs), UNets, and Fourier Neural Operators (FNOs) in order to better understand their strengths and limitations in modeling fracture. A further objective of this work is to evaluate the impact of three ensembling strategies (hard voting, soft voting and stacking) on the performance of these models, and to assess their effectiveness in improving predictive accuracy and robustness. This dataset extends our previous contributions to benchmark dataset development in mechanics, including the Mechanical MNIST dataset collection [31]. Specifically, it builds upon the Mechanical MNIST Crack Path dataset [49] by introducing a more advanced and more broadly relevant dataset. We refer to this new dataset, "PFM-Fracture" subset of our new "Nonlinear Mechanics Benchmark" dataset which consists of high-resolution finite element simulations of brittle fracture using PFM applied to domains with randomly initialized cracks.
Figure 1: Comparison between approaches to evaluating surrogate model efficacy: a) Standard examples from the literature such as tension, shear, and crack coalescences [46, 40], b) Our proposed benchmark dataset designed to provide a diverse set of challenges across initial conditions, load conditions, and energy decomposition methods.
The remainder of this paper is organized as follows. First in Section  2.1, we will start by discussing phase field modeling of brittle fracture. Next, in Section 2.2, we will present the specifications of the dataset that was created. After that, we will cover the details of the implemented baseline models and ensembling methods in Sections 2.3 and 2.4. Then, in Section 3, we will present the results of training them on the dataset as well as the outcome of applying ensembling methods to improve model predictions. We will compare the performance of these baseline models with each other in Section 3.5, and we will briefly go over the limitations of the metrics used in our training process and evaluation of machine learning models. Finally in Section 4, we present the conclusions and outlook of this work.
2 Methods
We begin in Section 2.1 by introducing the phase field fracture formulation for brittle materials and the three different energy decomposition methods implemented within our dataset. Then, in Section 2.2, we describe in detail the initial and boundary conditions used in our dataset, and the dataset file structure. Finally, in Section 2.3, we define the baseline machine learning models associated with this manuscript. We also note briefly that all code associated with the generation of this dataset and the implementation of these baseline models is available on GitHub, with a link provided in Section 5.
2.1 Phase Field Model Preliminaries
2.1.1 Phase Field Modeling of Fracture
Fracture is a common failure mechanism for engineering materials and structures [50]. From both a theoretical and computational standpoint, fracture remains one of the most challenging phenomena to model [43]. Early continuum mechanics approaches began with Griffith’s theory [51], which describes fracture as a competition between crack surface energy and elastic energy, and continued with the development of theory such as linear elastic fracture mechanics (LEFM) by Irwin [52], and the cohesive zone model (CZM) by Dugdale [53]. However, these methods rely on external criteria to predict crack nucleation and propagation, which prompted the search for alternative approaches, particularly approaches that would be amenable to simulation [43]. One promising and recently popular method with broad adoption in the literature is phase field modeling (PFM) of fracture [54]. Phase field modeling of brittle fracture offers a variational approach to modeling fracture and crack propagation by transforming the problem into the minimization of a potential energy function. This is achieved through regularization of the crack via a scalar field, called the phase field, using a characteristic length $l_{0}$ that defines the extent of regularization [44]. In the introduced phase field, the cracked regions smoothly transition from a scalar value of $\phi=1$ to the intact regions with a scalar value of $\phi=0$.
Figure 2: A general schematic of a mechanical system with a) the real sharp crack surface $\Gamma$ inside a domain $\Omega$ with boundary conditions, and b) the crack approximation via the scalar phase field with $\phi=1$ in the cracked region and $\phi=0$ in the intact region with a diffuse transition between the regions.
Miehe et al. [44] regularized the crack using a crack surface density function $\gamma(\phi,\nabla\phi)$. In this way, the original variational formulation of Griffith [51], which required the integration of crack surface energy over the unknown crack surface $\Gamma$, was transformed into an integration over the entire domain $\Omega\in R^{d}$ with Dirichlet boundary conditions on $\partial\Omega_{g}\subseteq\partial\Omega$ and Neumann boundary conditions on $\partial\Omega_{h}\subseteq\partial\Omega$ as shown in Fig. 2. With this definition, we can define total energy potential $\mathcal{E}$ as
$\mathcal{E}(\mathbf{u},\phi)=\Psi_{s}+\Psi_{c}-\mathcal{P}=\int_{\Omega}\psi_{0}(\mathbf{\mathbf{\epsilon}}(\mathbf{u}))\mathrm{d}V+\int_{\Gamma}G_{c}\mathrm{d}A-\int_{\Omega}\mathbf{b}\cdot\mathbf{u}\mathrm{d}V+\int_{\partial\Omega}\mathbf{t}\cdot\mathbf{u}\mathrm{d}A$ (1)
where $\mathcal{E}$ is a function of the displacement $\mathbf{u}$, the phase field $\phi$, and the potential energy functional $\mathcal{P}$ of external forces is a function of distributed body force $\mathbf{b}$ and traction on the boundaries $\mathbf{t}$. The first term, $\Psi_{s}$, is defined as the strain energy potential and $\psi_{0}$ is the strain energy density function as a function of infinitesimal strain tensor $\mathbf{\epsilon}(x)$. We define $\mathbf{\epsilon}(x)$ as:
[…]
where $\nabla^{s}(\cdot)$ is the symmetric gradient operator. Assuming a linear elastic material, the strain energy density function is then defined as:
$\psi_{0}=\frac{1}{2}\lambda\operatorname{tr}^{2}(\mathbf{\epsilon})+\mu\operatorname{tr(\mathbf{\epsilon^{2}})}$ (3)
with $\lambda$ and $\mu$ as the Lamé constants. The dissipated surface energy is defined as $\Psi_{c}$. The surface energy term can be regularized as:
$\Psi_{c}=\int_{\Gamma}G_{c}\mathrm{d}A\approx\int_{\Omega}G_{c}\gamma(\phi,\nabla\phi)\mathrm{d}A$ (4)
where $G_{c}$ is the critical energy release rate of the material. The crack surface density function is defined as:
$\gamma(\phi,\nabla\phi)=\frac{1}{2}\left[\frac{1}{l_{0}}\phi^{2}+l_{0}(\nabla\phi\cdot\nabla\phi)\right]$ (5)
where $l_{0}$ is the characteristic length. Literature suggests that $l_{0}$ should be treated as a material property that governs the regularization of a sharp crack [55, 43]. As $l_{0}\rightarrow 0$ the solution converges to the original fracture problem; however, capturing a small characteristic length requires a highly refined mesh in the transition zone leading to a significant computational cost [41]. After introducing the regularization to the crack surface energy term, the bulk strain energy term is now defined as:
$\Psi_{s}(u,\phi)=\int_{\Omega}g(\phi)\psi_{0}(\mathbf{\epsilon}(u))\mathrm{d}V$ (6)
where $g(\phi)$ refers to the energetic degradation function. With this background, the total energy functional can be written as:
$\mathcal{E}(\mathbf{u},\phi)=\int_{\Omega}g(\phi)\psi_{0}(\mathbf{\epsilon}(\mathbf{u}))\,\mathrm{d}V+\int_{\Omega}G_{c}\frac{1}{2}\left[\frac{1}{l_{0}}\phi^{2}+l_{0}(\nabla\phi\cdot\nabla\phi)\right]\,\mathrm{d}V-\int_{\Omega}\mathbf{b}\cdot\mathbf{u}\mathrm{d}V+\int_{\partial\Omega}\mathbf{t}\cdot\mathbf{u}\mathrm{d}A\,.$ (7)
Various variational and non-variational approaches have been proposed in the literature. In variational formulations, the stress field and phase-field evolution equations can be derived from Eq.7, as seen in [56, 44]. However, these formulations lead to nonlinear equilibrium equations, which are computationally expensive to solve. To address this issue, hybrid formulations have been introduced [45, 57], incorporating a linear displacement subproblem to reduce the computational cost. Given the complexity of each simulation sample—particularly due to the need to resolve crack coalescence and the irregularity of initial conditions—we have implemented the hybrid phase field formulation proposed by Ambati et al. [45], which maintains a linear momentum equation to reduce computational costs. With this approach, the linear momentum equation is written as follows:
[…]
where the history field variable $\mathcal{H}^{+}$ is adopted from [44] to enforce the irreversibility of the crack phase field evolution where $\psi_{0}^{+}$ is the tensile component of elastic energy density. Each of the displacement field and phase-field functional are minimized iteratively in an alternating minimization framework, which offers a robust but slow convergence [43]. An important feature of Eq.7 is that it is convex with respect to $\mathbf{u}$ and $\phi$, but is not convex with respect to the both [41]. This further necessitates the use of an alternating minimization method for solving the PDEs [43]. To enforce tensile-compressive asymmetry in the fracture response, the elastic energy density function must be decomposed into tensile and compressive parts where only the positive component should be degraded using the energy degradation function $g(\phi)=(1-\phi)^{2}$ as in Eq.9. With this approach, $\psi_{0}(\mathbf{\epsilon}(\mathbf{u}))$ is written as:
$\psi_{0}(\mathbf{\epsilon}(\mathbf{u}))=g(\phi)\psi_{0}^{+}(\mathbf{\epsilon}(\mathbf{u}))+\psi_{0}^{-}(\mathbf{\epsilon}(\mathbf{u})).$ (9)
In this work, we implement three different energy decompositions methods: volumetric-deviatoric decomposition proposed by Amor et al. [55], spectral decomposition proposed by Miehe et al. [44], and the star-convex decomposition proposed by Vicentini et al. [58]. Each of these decompositions will be discussed further in Section 2.1.2. We also briefly note that, though we do not cover them here, there have been several different approaches in the literature including non-variational phase field modeling that will allow for crack healing and nucleation under compression, and better prediction of crack nucleation in the absence of large pre-existing cracks by utilizing the material strength surface instead of energy splitting methods [59, 60, 61, 62].
2.1.2 Energy Decomposition Methods
In order to best represent the state of the art in phase field fracture simulations, our dataset contains simulation results from three different energy decomposition methods. The first two, spectral decomposition [44] and volumetric-deviatoric decomposition [55], have a long history of adoption by the community. However, multiple different alternative approaches to energy decomposition have been suggested in the literature [63, 64, 65, 66]. Recently, Vicentini et al. [58] offered a modification to the volumetric-deviatoric decomposition utilizing a parameter $\gamma^{*}$ to address the shortcomings of other decomposition methods in modeling crack nucleation and propagation. We provide an implementation of this method, the star-convex method, as an additional example of a simple tunable method allowing the calibration of tensile, compressive, and shear strength of a material under multi-axial loading.
Spectral Decomposition
An energy decomposition method, suggested by Miehe et al. [44], decomposes the strain into tensile and compressive parts by using the eigenvalues and eigenvectors of the strain tensor written as:
$\mathbf{\epsilon}^{+}$ $=\sum_{i}\langle\mathbf{\epsilon}_{i}\rangle_{+}e_{i}\otimes e_{i}$ (10)
$\mathbf{\epsilon}^{-}$ $=\sum_{i}\langle\mathbf{\epsilon}_{i}\rangle_{-}e_{i}\otimes e_{i}$ (11)
where $\mathbf{\epsilon}_{i}$ is the eigenvalue, $e_{i}$ is the corresponding eigenvector, and $\otimes$ refers to the dyadic product. This decomposition leads to the expression:
$\psi_{0}^{+}(\mathbf{\epsilon})$ $=\frac{1}{2}\lambda_{0}\langle\operatorname{tr}(\mathbf{\epsilon})\rangle^{2}+\mu_{0}\mathbf{\epsilon}^{+}:\mathbf{\epsilon}^{+}\,$ (12)
$\psi_{0}^{-}(\mathbf{\epsilon})$ $=\frac{1}{2}\lambda_{0}\langle-\operatorname{tr}(\mathbf{\epsilon})\rangle^{2}+\mu_{0}\mathbf{\epsilon}^{-}:\mathbf{\epsilon}^{-}$ (13)
Though this decomposition is computationally more intensive, it better models compression-dominated loadings, where all three principal strains are negative when compared to the volumetric-deviatoric decomposition approach [67, 43].
Volumetric-Deviatoric Decomposition
This decomposition method introduced by Amor et al. [55] decomposes the strain tensor into a deviatoric $\mathbf{\epsilon}_{D}$ term and a volumetric $\mathbf{\epsilon}_{V}$ term as follows:
$\mathbf{\epsilon}$ $=\mathbf{\epsilon}_{V}+\mathbf{\epsilon}_{D}$ (14)
$\mathbf{\epsilon}_{V}$ $=\frac{1}{3}\operatorname{tr}(\mathbf{\epsilon})\mathbf{1}$ (15)
$\mathbf{\epsilon}_{D}$ $=\mathbf{\epsilon}-\frac{1}{3}\operatorname{tr}(\mathbf{\epsilon})\mathbf{1}$ (16)
The resulting strain energy density potential is defined as:
$\psi_{0}^{+}(\mathbf{\epsilon})$ $=\frac{1}{2}\kappa_{0}\langle\operatorname{tr}(\mathbf{\epsilon})\rangle^{2}+\mu_{0}\mathbf{\epsilon}_{D}:\mathbf{\epsilon}_{D}$ (17)
$\psi_{0}^{-}(\mathbf{\epsilon})$ $=\frac{1}{2}\kappa_{0}\langle-\operatorname{tr}(\mathbf{\epsilon})\rangle^{2}$ (18)
[…]
Star-Convex Decomposition
We now introduce the recently proposed star-convex energy decomposition method, which can be viewed as a modification of the volumetric-deviatoric approach aimed at enhancing its flexibility by allowing for calibration of compressive $\sigma_{e}^{-}$ and tensile strength $\sigma_{e}^{+}$. In this approach, a $\gamma^{*}$ term is introduced to allow for a residual stress related to volume contraction $\langle-\operatorname{tr}(\mathbf{\epsilon})\rangle$, and the strain energy is then written as:
$\psi_{0}^{+}(\mathbf{\epsilon})=\frac{1}{2}\kappa_{0}\left(\langle\operatorname{tr}(\mathbf{\epsilon})\rangle^{2}-\gamma^{\star}\langle-\operatorname{tr}(\mathbf{\epsilon})\rangle^{2}\right)+\mu_{0}\mathbf{\epsilon}_{D}:\mathbf{\epsilon}_{D}$ (19)
$\psi_{0}^{-}(\mathbf{\epsilon})=(1+\gamma^{\star})\frac{1}{2}\kappa_{0}\langle-\operatorname{tr}(\mathbf{\epsilon})\rangle^{2}$ (20)
where $\gamma^{*}=\sigma_{e}^{-}/\sigma_{e}^{+}$. We note briefly that in this decomposition method, the original volumetric-deviatoric decomposition can be reconstructed by setting the term $\gamma^{*}=0$. In this work, we select $\gamma^{*}=5$ to match the original reference.
Figure 3: Geometrical description and boundary conditions of the three benchmark cases and the resulting crack path and force displacement curves to validate the implemented phase-field model. Here we show: a) the tension case, b) the shear case and c) the crack coalescence case under uniaxial tension. The crack patterns for the spectral decomposition is depicted in (ii). The force displacement curve for all three of our energy decomposition approaches are reported in (iii) alongside results from the literature.
2.1.3 Benchmark Validation Problems
When introducing a new phase field formulation and validating a phase field model implementation, it is standard to report performance on multiple benchmark examples [45, 68, 69]. Many different benchmark cases have been suggested in the phase field fracture literature, some with the aim of examining the effect of different modeling assumptions, and others with the aim of ensuring correct model implementation. At present, there are published benchmarks relevant to quasi-static fracture [45], dynamic fracture [70], hydraulic fracture [71, 72], crack coalescence [73], multi-axial loading, tensile loading, compressive loading, and shear loading [74].
However, despite the fact that comparing to standard benchmarks is common practice, in our analysis of the literature we found that it was often quite challenging to reproduce the benchmark results reported by others. Critically, phase field fracture approaches are sensitive to multiple modeling parameters, including basic mesh structure, characteristic length, convergence criteria, and the strategy for modeling pre-exsiting cracks [41, 43]. And, because sharing software and implementation code are not standard practice, it is especially difficult to isolate sources of discrepancy. To this end, we make a note of our implementation choices in Section 2.2, and share all of the code underlying our implementation of the models shown in this Section. To validate our implementation of the three approaches outlined in Section 2.1.2 under different types of loading and boundary conditions, we compare their performance to each other and to the literature on three standard benchmark problems. These popular benchmark cases include: (1) Tension loading on a single-edged notched specimen – referred to as the “tension case,” (2) shear loading on a single-edged notched specimen – referred to as the “shear case,” and (3) uniaxial tension loading on a domain with internal flaws where crack coalescence is expected – referred to as the “crack coalescence case.” We show our results in Fig. 3 and provide additional details below. As briefly mentioned in Section 2.1.1, an extremely fine mesh is required in the crack transition zones to accurately compute the crack surface energy. The literature suggests that the mesh size in the cracked region should be at least $h=l_{0}/2$ as suggested in [70, 44]. Another important aspect in implementing PFM models is the definition of initial pre-existing cracks. Consistent with previous reports from the literature [43], different methods for defining the initial crack can result in a different force displacement response, which may explain small discrepancies between our force-displacement curve results and the results from other implementations. In this work we have used the initial strain history defined in [70]
$\mathcal{H}_{0}(x)=\frac{\phi_{c}}{1-\phi_{c}}\begin{cases}\frac{G_{c}}{2l_{0}}\left(1-\frac{2d(x)}{l_{0}}\right)&\text{if }d(x)\leq l_{0}/2,\\ 0&\text{if }d(x)>l_{0}/2.\end{cases}$ (22)
Here, $\phi_{c}=0.999$ represents the phase-field scalar value assigned to pre-existing cracks, while $d(x)$ denotes the shortest distance between points within the domain and the crack.
Figure 4: Visualization of crack branching near the boundaries for volumetric-deviatoric and star-convex decomposition. Note the qualitative differences between the three approaches.
Tension case
The tension case domain is illustrated in Fig. 3 a-i. Specifically, it is a $1\times 1$ mm square containing an initial crack with length 0.5 mm in the middle of the left edge. The top boundary experiences a uniform vertical displacement with a rate of $\Delta u=1E-5$ mm for the first 500 steps, and $\Delta u=1E-6$ mm for the next 1500 steps. For this case, we use an unstructured mesh with linear triangular elements refined in the central band of the domain where the crack is expected to propagate. The smallest element size inside the region is $h=0.001$ mm, and the largest element size outside the crack zone is $h=0.01$ mm. The convergence criteria tolerance value was set to $41E-$ 4. The initial crack was modeled by using an initial strain history field following Borden et al. [70]. All other parameters were set to match the recent literature [45] with Young’s modulus set to $E=210$ GPa, $\nu=0.3$, $G_{c}=2.7$ N/mm and characteristic length $l_{0}=4E-3$ mm. As shown in Fig. 3 a-ii, where the typical resulting crack pattern is illustrated, and in Fig. 3 a-iii where the force displacement curves are illustrated, the results obtained from our simulations for the different energy decompositions are in good agreement with the literature [45].
Figure 5: Overview of our dataset: a) Schematic showing the boundary conditions and the geometric properties of the domain, with $n$ number of cracks, where the $\theta$ orientation and center coordinate $(x,y)$ of each crack is randomly sampled for each simulation case. The crack length $l=0.25$ mm is fixed in all simulations, b) For each simulation, phase field and x and y components of the displacement in multiple loading steps and the boundary force vs. boundary displacement curves are saved.
Shear case
For the shear case, all physical parameters (Young’s modulus $E$, critical energy release rate $G_{c}$, characteristic length $l_{0}$ and the poisson’s ratio $\nu$) match the tension case. We use a uniform structured mesh with $256\times 256$ Q1 elements, following the original implementation by [45]. The top boundary is subject to shear with a constant rate of $\Delta u=1E-5$ mm. The boundary conditions are shown in Fig. 3 b-i , a representative resulting crack pattern is shown in Fig. 3 b-ii, and the load displacement curves are shown in Fig. 3 b-iii. Consistent with the literature [45], there are differences in the force displacement curve response using volumetric-deviatoric decomposition when compared to the other models. The spectral decomposition and the star-convex decomposition approaches show a similar response.
Crack coalescence case
Beyond the standard tension and shear benchmark cases, we also implemented a benchmark case from Zhou et al. [73] due to its similarity to the final dataset proposed in this work. In Zhou et al., initial flaws are modeled as geometry-induced cracks inside the mesh, but in our model we use a strain history field for defining the initial cracks. To closely match the force-displacement curve presented in Zhou et al. [73], we used a characteristic length of $l_{0}=0.12$ mm and a regionally refined unstructured mesh with triangular elements where the smallest element size within the region is $h=0.05$ mm and the largest mesh element size is $h=0.25$mm. As illustrated in Fig. 3 c-i, the domain is under uniaxial tension. Young’s modulus is set to $E=30$ GPa, $\nu=0.333$ and $G_{c}=3e-3$ N/mm. As can be seen in Fig. 3 c-ii-iii, the crack pattern and the force displacement curves are in good agreement with Zhou et al. [73] which uses the spectral decomposition approach. However, it is worth pointing out that in this case, using either the volumetric-deviatoric or the star-convex decomposition results in branching of the crack pattern near the boundaries as shown in the Fig. 4.
In our work, all simulations are implemented using the opensource finite element package FEniCSx [75, 76, 77, 78].
Figure 6: Figure showing the initial crack pattern and final crack pattern for three samples of each dataset subcategory. From these illustrative results we see that in some cases almost no difference between the three energy decomposition methods (see $3^{rd}$ row of bi-axial tension case and $2^{nd}$ row of shear case), moderate difference ($1^{st}$ row of bi-axial tension case and shear case) or extensive difference ($3^{rd}$ row of shear case) in the resulting final crack pattern.
2.2 Dataset specifications
The version of our dataset published in conjunction with this paper contains 6,000 two dimensional quasi static simulations of phase field modeling for brittle fracture. We simulate domains with internal initial cracks under biaxial tension and shear boundary conditions. For the shear case, the top and bottom boundaries are also subject to $\phi=0$ Dirichlet boundary condition as in many examples without this condition, the fracture initiated and propagated along these boundaries, resulting in uninformative fracture patterns that are not useful for training ML models. We run all three energy decomposition methods (1000 initial conditions $\times$ 2 boundary conditions $\times$ 3 energy decomposition methods $=$ 6000 simulations in total). Schematics of representative geometries and applied boundary condition are shown in Fig. 5a. The overall dataset structure is as follows. First, each simulation is categorized under one of 6 sub-categories named as case-energy_decomp for example: tension-spect. Each sample is saved under the random seed that was used for generating the random internal cracks in order to enhance reproducibility. The random parameters in each sample are the number of internal cracks which is an integer between 10-20, and the coordinates and orientations of the crack inside the domain. All of the cracks have the same length $l=0.25$ mm. For each sample, at 100 simulation steps, we save the phase field, the x and y components of the displacement field on the mesh, the force displacement data on the support edges, and the data is saved in HDF5 format. The simulation parameters including Young’s modulus $E$, Poisson’s ratio $\nu$, critical energy release rate $G_{c}$, characteristic length $l_{0}$ and the mesh resolution are listed in Table 1. A uniform structured mesh with bi-linear elements were used with 4 quadrature points in each element.
Table 1: Material parameters and mesh resolution used in the simulations. Please note that the code to reproduce these data is also freely available, see Section 5.
Parameter Value
$E$ (Young’s modulus) 1000 GPa
$\nu$ (Poisson’s ratio) 0.3
$G_{c}$ (Fracture energy) 1 N/mm
$l_{c}$ (Length scale) 0.01 mm
Mesh resolution $800\times 800$ Q1
The dataset can be accessed in both a $128\times 128$ grid format, and an original mesh resolution format. The $128\times 128$ format offers a good balance of information and data size maintaining a standard resolution for ML applications and ease of use, while the original mesh resolution format enables flexibility. In Fig. 6, we also show that different energy decompositions sometimes – but not always – result in significantly different outcomes.
2.3 Baseline Models
In this Section, we briefly summarize the baseline models implemented on these data. Following the example set by PDEBench [35], we choose three standard baselines: An extension of Physics Informed Neural Networks (PINNs) called Deep Ritz Method (DRM) [79], Fourier Neural Operators (FNOs), and U-Net convolutional neural networks (UNet). Standard “out-of-the-box” implementations of these baseline models are not designed with our phase field fracture dataset in mind. Thus, in all three cases, we elected to iterate on architecture and hyperparameters up to a point. Specifically, our goal was to make a balanced investment in baseline model selection – sufficient effort to generate meaningful results, while leaving advanced machine learning model development in this area to future work by both us and others. In addition, we made the pragmatic choice of prioritizing model implementations where the code was published alongside the paper. All the models were implemented using PyTorch [80] and the process of logging and hyperparameter tuning was done using the Weights and Biases tool [81]. After preliminary investigation, we selected a recently published PINN formulation designed specifically for phase field fracture problems [40], the “out-of-the-box” implementation of FNOs from PDEBench [35], and the UNet architecture used in our previous work on calibration for classification problems in mechanics [82]. Further details follow.
2.3.1 Physics Informed Neural Networks
The goal of training a Physics-Informed Neural Network (PINN) is to approximate the solution to a partial differential equation (PDE). The concept was introduced in [7], and since then there has been an explosion of related work in the literature [83, 84, 85, 86, 87]. In some cases, neural networks have been used to minimize the residual of the strong form of the PDE, as in [7], while in others, the variational form of the PDE has been used, as in [85, 88, 79]. There have been multiple recent attempts to apply PINNs to phase-field models of brittle fracture, for example [46, 47, 89, 40, 39, 48, 90]. In some of these studies, to mitigate the computational cost of training a PINN for every loading step, different transfer learning approaches have been explored. For instance, the trained weights from one loading step can be used as a good initialization point for the next step [40], or the weights of all layers except the last can be kept fixed after training during the first step, and then only the weights of the last layer of the feedforward network are retrained [39]. A typical PINN architecture consists of multilayer perceptrons (MLPs), and usually weighted combinations of multiple loss terms that are mainly physics-informed as a form of a PDE residual and can also have supervised data-driven terms, on a set of collocation points inside the domain or on the boundary [7]. PINN based models are usually trained for a great number of iterations to drive the residual toward zero.
For our representative baseline model, we explore the process of minimizing the PINN loss function for only a single set of boundary conditions and initial cracks over a predefined number of loading steps. In our preliminary explorations (not reported in detail in this paper), we attempted both the vanilla PINN formulation using the DeepXDE Python package [91], and the DRM model developed by Manav et.al.  [40]. The DRM uses the quadrature points of the mesh as the collocation points, and is formulated to minimize the energy functional of the variational problem [46, 40]. For the main PINN baseline model presented in this paper, we have used the PINN architecture and code provided by Manav et al. [40] as a representative example from the recent literature of a PINN formulation with phase field fracture as the target application. We chose this approach because the implementation was well documented and openly available and was benchmarked on the crack coalescence case which aligned closely with our dataset. To make a fair comparison, a uniform unstructured mesh with the same mesh size as the FEA solution described in Section 2.2 with triangular elements was provided as the input to the model. After experimenting with different hyperparameters such as width and depth of the hidden layers, we settled on 6 hidden layers, each having 100 neurons and with characteristic length set to 0.01 mainly because other configurations were computationally prohibitive. As stated in the original paper, training a model can take up to an order of magnitude longer for much simpler crack pattern cases. We note briefly that though this method has demonstrated efficacy on simpler initial crack patterns (e.g., single-edge notched specimen under tension and shear and crack coalescence), it has not been rigorously tested on more complex patters such as the ones provided in our dataset.
2.3.2 Fourier Neural Operator (FNO)
Fourier Neural Operators were proposed for operator learning, mapping infinite-dimensional input functions to output functions [92]. Fourier Neural Operators (FNOs) are among a plethora of operator learning approaches that parametrize the integral kernel in the Fourier space using neural networks [93, 94, 95].
In this work, we have followed the PDEBench [35] implementation and used the same architecture and hyperparameters as the original paper by Li et al. [92]. Specifically, this implementation has four Fourier layers, with the modes in each direction set to 12 with Gaussian Error Linear Unit (GELU) activation functions. We trained the model for 300 epochs with the mean squared error (MSE) loss and the Adam optimizer [96] setting the initial learning rate to 0.001 with a scheduler set to halve the learning rate every 100 epochs, and a batch size of 4. The FNO model was trained autoregressively using three channels for the input (phase field, displacement field in x direction, and displacement field in y direction) with ten steps for autoregressive rollout for five different seeds. In addition, we 0-1 normalized the input and output of the displacement fields of the FNO for each loading step based on the applied displacement on the boundaries.
2.3.3 UNet
Here we use the UNet architecture implemented in [82]. Due to their ease of implementation and popularity in the broader literature [35, 34], UNets are a natural choice as a baseline model [97]. After preliminary exploration of multiple potential UNet architectures and loss functions, we settled on the UNet architecture used in previous work by our group for investigating model calibration for classification [82]. In contrast to our FNO implementation, we did not train the UNets autoregressively because in our initial attempts, autoregressive training led to unstable behavior. And, following preliminary exploration, we only trained the UNets with the phase field channel. Because the UNet struggled with the limited number of training samples, we applied data augmentation such as random rotations and horizontal and vertical flips on the fly to improve its performance. In contrast, the FNOs did not require such augmentation, since they were trained autoregressively over the time steps and did not suffer from the same limitation. We set the batch size to 16 for 100 epochs using the Adam optimizer with the learning rate set to 0.0001. For the UNet we used the sum of Dice and Focal loss with $\gamma=2$ and $\alpha=0.8$ [98] with equal weights and we used 0.4 threshold for binarizing the ground truth. The same dataset and data split was used for the UNet as the FNO.
2.4 Ensembling methods
We trained both the FNO and UNet models with five different random seeds. To showcase performance in this paper, we used the dataset with bi-axial tension boundary condition and the spectral decomposition energy split with a $128\times 128$ input resolution. The dataset was split into $80\%$ training data, with the remaining $20\%$ split into $60\%$ as the meta train split to train the stacking ensemble, $20\%$ as the meta validation set for tuning the hyperparameters of the stacking model, and the final $20\%$ used as the meta test set. As seen in Fig. 10, models trained with different random seeds often predict different results. Compellingly, and consistent with our previous work in the area [82], the multiple different results often look “plausible” which motivates the general idea of exploring methods that can combine these different results with an eye towards future exploration from an uncertainty quantification perspective [99]. In line with the goals of this paper, we provide baseline results from performing straightforward ensemble averaging across the five model seeds using majority and unweighted averaging also known as hard and soft voting respectively [100, 101] and stacking methods [102]. A brief explanation of each approach follows.
Hard voting is a well known ensembling method also referred to as “majority voting.” In this simple approach, we aggregate the class label predictions from different models, and the majority vote determines the final prediction. Soft voting is another well known and straightforward ensembling method that averages the class label probabilities across different models [100]. It has been observed that even using these simple ensembling methods can improve the performance of the models such as UNet [82], and different random initializations of the models can lead to learning different patterns and modes of the data and ensembling the results of them would reduce the variance of the final prediction [103, 5, 101]. Finally, stacking, originally introduced in [102], is an ensembling method that trains a meta-learner on the predictions of the base models to learn a linear or nonlinear combination of their outputs, producing a final prediction for the class label. To perform our baseline model of stacking, we used a simple architecture with 81 parameters that is comprised of two 2D convolutional layer with kernel size 1 and 3 to denoise and capture local features on all five channels and then a ReLU activation function and a final 2D convolutional layer that aggregates all five channels into the final output. We have observed that in this small network the initialization is very important to get a good result. As the probability of having a cracked pixel inside the domain is very low (approximately $2\%$) we initialized the weights of the first and the last layer using Xavier initialization [29]. Additionally we initialize the bias of the first layer to 0.2, enabling the model to start from a soft voting state in which each of the five sub models is given equal weights. The $3\times 3$ depthwise layer is initialized with the identity stencil with no bias. The reasoning behind this initialization was so that the initial pass through the network would start from a stable point and to prevent dead ReLU activations. We trained the stacker model for 50 epochs on the meta train split of both UNet and FNO predictions. As the stacking method in its most basic form can approximate soft voting by learning equal weights for each channel, it should be noted that its true potential lies in extending it with added contextual windows and nonlinearities. This allows the network to learn a nonlinear combination of the sub models’ outputs and to learn from weaknesses and strengths of them across different regions of the field. Though many more advanced ensembling techniques exist [104, 105, 106] and are considered “state of the art,” we consider them beyond the scope of this paper where our goal is to provide a set of baseline models.
Figure 7: Figure showing: a) The initial condition and the resulting FEA simulation under bi-axial tensile loading and b) The results of training the PINN model (12 separate models) with three energy decomposition methods across four different seeds. We note that the individual PINN models are converging to different crack patterns. Though some of these models show plausible results, none were able to predict the correct crack pattern.
3 Results and Discussion
In this Section we investigate the results of training the baseline models and discuss approaches to improve and evaluate their performance. Sections 3.1 to 3.3 present the performance, limitation and potential future directions for PINNs, FNOs and UNets. Then in Section 3.4, we investigate the improvement as a result of using different ensembling approaches. In Section 3.5, we provide a comparison of the baseline models, and finally in Section 3.6 we briefly go over the implications of using different error metrics for ML model training and evaluation.
Figure 8: This figure presents the predicted phase field and displacement fields for each FNO model that was trained on the dataset. It shows the results for a representative sample at the $50^{th}$ percentile of performance with a Dice score of 0.6869. The top row shows the ground truth and each subsequent row shows the results of each FNO seed. Additional examples are shown in Fig. 13
3.1 Physics Informed Neural Network Baseline
To investigate the ability of PINNs to predict phase field fracture behavior, we began with a vanilla implementation using the DeepXDE package [91]. Consistent with previous studies [40, 48], we found that using the strong form of the PDE only produces meaningful phase field fracture results in the 1D setting. Thus, for the baseline PINN model shown here, we adopted the Deep Ritz method (DRM), proposed by Manav et al. [40]. Specifically, we separately trained 12 models under the bi-axial tension boundary condition. We investigated all three energy decomposition methods (spectral, volumetric-deviatoric, and star-convex) with four different seeds each. In the FEA setting, these simulations were displacement controlled, and all three energy decomposition strategies resulted in the same crack pattern. In Fig. 7a, both the simulation initial condition and the ground truth FEA output (identical for all energy decomposition methods) are shown.
The results of training the DRM model for all four seeds are shown in Fig. 7b. As can be seen from the illustrations, none of the models converged to the correct solution, and each of them resulted in a distinctly different crack pattern. In the original paper proposing the DRM model [40], there was a crack coalescence test that showed good agreement with the ground truth finite element solution. However, we note that: (1) the reported PINN crack pattern differs slightly from the finite element derived ground truth, and (2) it is a test with only three initial cracks and is therefore much simpler than the example investigated in Fig. 7. With only three initial cracks (all oriented in the same direction and with even spacing, see Fig.17 in [40])the inaccuracies in approximating the energy landscape are less probed than in the example shown in our paper where errors manifest more clearly in the predicted crack paths.
We note briefly that in designing this study we chose a uniform mesh with the same mesh size as the FEA simulations to train the PINNs ($0.0025$ mm). Due to the high computational cost of training each PINN model (approximately 24 hours to simulate up to the same loading step as the FEA model on one NVIDIA V100 GPU), we did not iterate significantly to tune the network architecture or the number of trainable parameters of the model. Critically, we also note that poorly performing examples may still lead to visually convincing but physically incorrect results. This highlights the importance of incorporating additional physics-based constraints and error evaluation approaches, as these misleading results during test-time inference can give a false sense of accuracy and reliability. Exploring such modifications is one promising direction for improving PINN performance in the future, and we hope that our dataset will encourage further development in the area.
Figure 9: This figure presents: a) The histogram of the Dice score distribution of the FNO models trained with five different seeds and ensembling methods, and b) representative test samples showing the performance of each ensembling method. For each sample, the initial condition is shown in the first column and the corresponding ground truth contour is shown on the second column. The subsequent columns display the result of ensembling methods with the ground truth contour overlaid in red for visual comparison.
Figure 10: This figure presents: a) The histogram of the Dice score distribution of the UNet models trained with five different seeds and ensembling methods, and b) representative test samples showing the performance of each ensembling method. For each sample, the initial condition is shown in the first column and the corresponding ground truth contour is shown on the second column. The subsequent columns display the result of ensembling methods with the ground truth contour overlaid in red for visual comparison.
3.2 Fourier Neural Operator Baseline
Unlike the PINNs approach described in Section 3.1, we were able to achieve reasonable baseline model results with the standard FNO model described in Section 2.3.2. Thus, in this Section, we present both quantitative metrics and qualitative outputs for trained FNO model performance. In Fig. 9a-b, we show the quantitative and qualitative performance of the FNO model for all five individual seeds, and the ensemble results across all five seeds. The ensemble results will be discussed in more detail in Section 3.4.
In Fig. 9a, we plot a histogram of the Dice score across all 40 test samples evaluated with the trained FNO models. As indicated by the vertical line labeled “Dice Mean Seed 1-5” in the Fig. 9a histogram, the average Dice score across the test set and across all seeds is 0.6208. We can see in Fig. 8 that FNO models predicted a visually less well-defined and smoother phase field solution when compared to the FEA Ground Truth. A plausible explanation for this outcome is related to a well-known limitation of neural networks referred to as spectral bias, which indicates the preference of neural networks to learn low frequency features at the expense of high frequency features [23, 107, 108, 24]. Another interesting observation from Fig. 8 is that although the predicted displacement fields may deviate from the ground truth, they remain consistent with their corresponding phase fields. This suggests that the models have captured some underlying physical relationships through training on this dataset.
Looking ahead, there are several promising directions to improve upon these baseline models. For example exploration of various data augmentation strategies, ranging from traditional techniques to more advanced generative AI based approaches. application of different data augmentation techniques from simpler methods to generative AI based methods. As shown in Fig. 9b, which presents crack pattern predictions ranked from best to worst based on the soft voting method (discussed further in Section 2.4), we observe that examples with lower Dice scores typically have large quantitative and qualitative deviations from the ground truth. Thus, although the top-performing predictions (e.g., those in the upper row of Fig. 9b) show strong agreement with the FEA ground truth, the substantial discrepancies in lower-performing cases highlight significant room for improvement.
3.3 UNet Baseline
In this Section, we evaluate the performance of the UNet models on the same dataset to assess predictive accuracy across different random seeds and ensembling strategies. As outlined in Section 2.3.3, the standard UNet architecture without any hyperparameter tuning was trained on the bi-axial tension dataset with spectral energy decomposition, with five different seeds.
Fig. 10a shows the histogram of the Dice score distribution across all 40 test samples evaluated with the trained UNet models and ensembling approaches (discussed in Section 2.4). The UNet models on average scored 0.6804 in Dice score among different seeds. Fig. 10 also highlights that the models – likely by design as they are trained on the binary phase field – produced sharper predictions. Note that unlike the FNO models, the displacement fields were not used during UNet training.
Figure 11: The Figure shows a heatmap of the performance of each ensembling method on the predictions of a) UNet and b) FNO submodels based on the chosen threshold that was used to binarize the prediction and the corresponding ground truth to calculate the Dice score.
3.4 Ensemble approaches
In addition to evaluating individual models, we investigated the impact of the three ensembling methods described in Section 2.4 (hard voting, soft voting, and stacking) on overall prediction performance. To compute the Dice score, which requires a binarized prediction field, we were required to threshold both the model prediction fields and the ground truth fields to delineate between “undamaged” (i.e., less than threshold) and “damaged” (i.e., more than threshold) regions. Because test performance is dependent on this choice of threshold, we performed a threshold search over the range [0.1, 0.6] using the validation set for both the FNO and UNet models separately. The reason for choosing this range was because the models, particularly the FNOs, tended to produce values in range 0.1 to 0.6, as illustrated in Fig. 8. Thus, applying a higher threshold (e.g. close to 0.9) would result in omitting most of the predicted crack regions, as only a small fraction of the predictions would surpass the cutoff. This would lead to very low Dice scores, and would prohibit us from being able to effectively study and compare the performance of the baseline models. In Fig. 11a and b, we show the resulting heatmaps for selecting both a ground truth threshold and a prediction threshold for the purpose of scoring each ensembling method. In these plots, the red $\star$ indicates the best performing threshold combination on the validation set, which we selected and then subsequently evaluated on the test set.
Table 2: Dice scores of ensembling methods for FNO and UNet models.
Ensembling Method FNO UNet
Hard Voting 0.6831 0.7321
Soft Voting 0.6804 0.7315
Stacking 0.7333 0.7325
Mean Seeds 1-5 0.6208 0.6804
All ensemble results shown in Fig. 9b and Fig. 10b and listed in Table 2 reflect these threshold combinations. As widely observed in the literature [5, 82], ensembling improves the prediction performance of both model types. However, the magnitude of this improvement differs between the two. For the FNO predictions, Hard voting improved the performance of the model by $\approx 10\%$ to 0.6831, the soft voting by $\approx 6.8\%$ to 0.6804 and the stacking achieved the highest boost, improving the performance by $\approx 18.1\%$ to 0.7333. The stacking method with non linearity through ReLU activation function and context aware filters as described in Section 2.4 outperformed the other ensembling methods. One possible explanation for this outcome is that training a meta learner like the stacker model allows the network to act as a localized router learning which base model’s output to prioritize for different types of crack patterns. For example, one initial model might perform better on horizontal cracks while another may more accurately capture crack coalescence. In the future, we could also integrate this stacking model into the architecture itself to form a simple Mixture of Experts (MoE) model [109]. Looking at both Figs. 9 and 10, we can see that the contextual filters of the stacking model helped to filter out False positives and thus considerably improved the prediction accuracy.
In contrast, all three ensembling methods resulted in almost identical performance improvement when applied to the UNet. Hard voting improved the Dice score by $\approx 7.5\%$ to 0.7321, soft voting by $\approx 7.3\%$ to 0.7315 and stacking by $\approx 7.6\%$ to 0.7325. We anticipate that this difference is likely due to the lower number of false positives of the individual UNet predictions.
3.5 Overall Baseline Model Performance – Methods Comparison
Standard PINNs using the strong form of the PDE as the loss function only work on the 1D settings of phase field problems. Using the Deep Ritz Method (DRM) to incorporate the energy functional allows training in 2D. but it still fails to achieve accurate solutions in complex cases such as the samples in the current dataset. Moreover, the computational cost of this approach is extremely high (an order of magnitude more than FEA solutions) making it impractical for many applications.
When comparing the performance of the baseline FNOs to the baseline UNet, we can see that the FNOs produced more smoothed out predictions and struggled to accurately model high frequency features of the domain such as the crack paths. In contrast, the UNet models with Focal-Dice loss function, generally achieved higher Dice scores on the dataset compared to the FNO models and produced sharper predictions. The ensembling methods resulted in a bigger relative improvement on the FNO predictions of approximately $18\%$ when using the stacking method achieving a comparable performance to the UNet models with a Dice score of 0.7333. This highlights the potential of ensembling methods to mitigate some of the limitations of operator learning architectures especially in challenging prediction tasks.
That being said, there are several promising directions that could be explored to improve upon the baseline model performance and reliability from both the FNO and UNet starting points presented here. First, adopting more advanced UNet-based architectures such as ConvNet UNets [110] or Multires-WNet [111], which extract multi-scale features, could lead to higher performance on these data. Another promising avenue is to incorporate more robust autoregressive training strategies, such as push forward and temporal bundling techniques introduced by Brandstetter et al. [112], to stabilize training over time sequences. Lastly, increasing the number of retained Fourier modes in the FNO architecture may further improve its capacity to model high frequency features and help improve baseline performance. Overall, we anticipate significant future work in this direction. The purpose of this paper is to showcase a challenging problem, and provide a starting point upon which future improvements can be made.
3.6 A Note on Performance Metrics
In addition to potential performance enhancements via iterating on the ML modeling approach, we also note that for the phase field fracture problem specifically there are multiple directions for improvement beyond model selection. For example, in this study we can readily see the inherent limitations of typical performance metrics such as Mean Squared Error (MSE) – both as the loss function for the network, and as the overall performance evaluation metric. Critically, it has been observed in other studies that MSE tends to emphasize errors of the low frequency features, and encourage learning an averaging approach in the prediction fields [113, 111]. Different approaches have been taken to address the problem of retrieving the high frequency features of the data. These approaches can vary from using multi scale [114, 24] and multi resolution architectures [111], using Fourier based metrics [35], or Fourier based features [115].
In this study, we have addressed the limitation of MSE by using the Sørensen–Dice score for evaluating the performance on the phase field channel of the predictions – in the literature also known as the $F_{1}$ score. This metric is used very commonly in semantic segmentation and classification tasks. This score can be calculated as follows:
$DSC=\frac{2|\hat{Y}\cap Y|}{|\hat{Y}|+|Y|}=\frac{TP}{TP+\frac{1}{2}(TP+FP)}$ (23)
where the $\hat{Y}$ is the binarized prediction field and $Y$ is the binarized target field. Here we can also see the metric in the context of binary classification where TP stands for True Positive and FP stands for False Positive. The main motivation behind choosing this metric is that the crack patterns occupy a very small portion of the field (approximately $2\%$ of the pixels), with the majority of pixels belonging to the background. This creates a highly imbalanced dataset, making this metric well suited for evaluating performance in fields with sparse, intricate patterns where the true negative classes, the undamaged regions, occupy a large portion of the domain. As can be seen in equation 23, the Dice score is not affected by the true negatives. While MSE might be a suitable choice for smooth, dense and continuous by nature fields such as displacement fields providing a reliable measure of overall field accuracy, it tends to be dominated by undamaged or background regions of the domain in the phase field channel of the data, making it insensitive to small features of the domain.
While Dice score provides a useful measure for this case, it has notable limitations. It does not penalize the discontinuous crack paths, for example in Fig. 9 b all of the samples other than the best performing one shows visible discontinuities in the crack trajectory, and also can produce higher scores for samples with more initial cracks or those where the cracks travel a longer and more complex trajectory within the domain. This behavior can be seen in Fig. 9b when comparing the best and worst examples. Samples that have more initial cracks tend to get a higher score in comparison to those with a smaller number of initial cracks. These limitations highlight the need for more physically grounded or geometry aware evaluation metrics that account for topological correctness, connectivity, and adherence to physics based constraints. For example, future studies could explore alternative hybrid metrics that combine pixel overlap with physics informed loss components to better assess prediction quality in fracture simulations.
4 Conclusion
In this work, we implemented a hybrid phase field model for brittle fracture incorporating three different energy decomposition methods: spectral, volumetric-deviatoric, and star-convex. The model was verified against standard benchmark cases commonly used in the literature and implemented using the open source finite element software FEniCSx [75, 77, 76, 78]. Based on the verified implementation, we generated six datasets combining two boundary conditions, bi-axial tension and shear with the three aforementioned energy decompositions. We then trained three baseline models, PINN, FNO and UNet on these datasets to benchmark their performance. Our results indicate that PINN based approaches have limited applicability in this context, while UNet and FNO models are capable of producing more reasonable predictions. We also study three ensembling methods from simple to more complicated: hard voting, soft voting and stacking. We observe that ensembling methods can improve the performance of the models in general but their performance can differ based on the performance of each submodel. Finally we note the importance, and the effects of metrics used for evaluating and training these models and how they should be used to effectively be able to train and evaluate the performance of ML based solutions. We have published the dataset and open sourced the code to reproduce the dataset and the benchmark models to enable other research groups to build up on them to exceed the baseline performance.
Despite achieving reasonable results with two of the baseline models, we observed that these models are highly sensitive to initialization and are prone to converging to different local minima. This behavior reflects the inherent non convexity of the underlying optimization problem and suggests that our dataset can serve as a challenging benchmark for evaluating the robustness of different modeling approaches. In particular, it provides a valuable opportunity to study the effects of training hyperparameters, random seeds, loss functions and training strategies on model performance and convergence. We can also anticipate future studies in several directions, including model calibration, development of more effective ensembling strategies, design of better metrics and loss functions that can help with spectral bias in neural networks by helping to capture high frequency features such as crack paths and the effects of data augmentations based on generative AI models.
5 Additional Information
The full resolution and the 128x128 downsampled version of the PFM-Fracture dataset are publicly available on the Harvard dataverse at (https://dataverse.harvard.edu/dataverse/PFM-Fracture). Due to the large file size, the full resolution dataset is divided into two parts. For example the shear case with spectral energy decomposition is stored as shear-spect-fullres-part-1 and shear-spect-fullres-part-2. The corresponding downsampled version is named shear-spect-128. All code for reproducing the dataset is available on GitHub at (https://github.com/erfanhamdi/pfm_dataset) and the implementation of baseline models and ensembling methods can be found at (https://github.com/erfanhamdi/pfm_bench).
6 Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
7 Acknowledgments
This work was made possible with funding through the Boston University David R. Dalton Career Development Professorship, the Hariri Institute Junior Faculty Fellowship, the Haythornthwaite Foundation Research Initiation Grant, and the U.S. National Science Foundation Cyberinfrastructure for Sustained Scientific Innovation program, under the "Elements: Curating and Disseminating Solid Mechanics Based Benchmark Datasets" project CMMI-2310771 and the U.S. Department of Energy, Advanced Scientific Computing Research program, under the "Resolution-invariant deep learning for accelerated propagation of epistemic and aleatory uncertainty in multi-scale energy storage systems, and beyond" project (Project No. 81824). This support is gratefully acknowledged. We also acknowledge the support of Saeed Mohammadzadeh for his insightful comments on phase field fracture modeling with FEniCS and data curation, and Boston University’s Research Computing Services for providing computing resources.
Appendix A Mesh Refinement Study
In order to study the effect of the mesh size on the results of the PFM simulations, we perform a mesh refinement study on a single initial pattern for both boundary conditions and all three energy decomposition methods.
Figure 12: This figure presents the mesh study for a) bi-axial tension case and b) shear case with force displacement curves and crack patterns across mesh sizes from a quadrilateral $200\times 200$ mesh to $1000\times 1000$ mesh. As can be seen the difference between the $800\times 800$ mesh and $1000\times 1000$ mesh is negligible both in the force displacement curve and final crack pattern.
Appendix B Additional results of evaluating FNO model on the test dataset
In Section 3.2, we presented the performance of the FNO model on the sample with median Dice score, here we present four additional samples from the best performing to the worst performing sample.
Figure 13: This figure presents the predicted phase field and displacement fields for each FNO model that was trained on the dataset. From the best performing sample with Dice score of 0.8059 to the worst performing sample with Dice score of 0.5588.